OpenStudy (anonymous):

What is the value of p? x=1/4y^2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

You can't find the value of p with that information.

OpenStudy (anonymous):

haha maybe y or x??

OpenStudy (anonymous):

You need more information.

OpenStudy (mertsj):

Is this supposed to be x = 4py^2?

OpenStudy (mertsj):

So 4p = 1/4 and p = 1/16?

OpenStudy (anonymous):

man i think mertsj gets the mind reader award for this one!

OpenStudy (anonymous):

the quetion says "p" thats y im confused

OpenStudy (mathmate):

I think it's more like: y^2=4px or x^2=4py depending on the orientation http://en.wikipedia.org/wiki/Parabola

OpenStudy (mertsj):

Except I think it's supposed to be x^2=4py so I think we have to rearrange it.

OpenStudy (anonymous):

the equation is x+1/4y^2

OpenStudy (mertsj):

So y^2=4x and 4p=4 and p = 1

OpenStudy (anonymous):

THANK YOU!!!! :)

OpenStudy (anonymous):

I understand it now

OpenStudy (mertsj):

yw